# Part I: General framework, application to Hartree-Fock and DFT

### v1

@@ 3 Dynamics on Grassmann manifolds (FCI, TDHF, TDDFT)
where $r,n$ are positive integers such that $11\leq r\leq n-$ 1 (the trivial cases $\mathcal{M}={\rm Gr}_{\mathbb{C}}(0,n)=\{0\}$ and $\mathcal{M}={\rm Gr}_{\mathbb{C}}(n,n)=\{I_{n}\}$ are uninteresting). Such manifolds appear in quantum dynamics in at least two settings:
@@
    pure-state dynamics of generic isolated finite-dimensional quantum systems described by the linear Schrödinger equation, or equivalently the linear quantum Liouville equation (see Section 3.3). In computational quantum chemistry, this corresponds to the FCI level of theory;
@@
    mean-field models for many-body fermionic systems. This framework encompasses the cases of both TDHF and adiabatic TDDFT, for which the dynamics is governed by a system of coupled nonlinear Schrödinger equations, or equivalently a nonlinear quantum Liouville equation (see Section 3.4).
Before elaborating on these two fundamental examples, the geometry and Kähler structure of the complex Grassmann manifold ${\rm Gr}_{\mathbb{C}}(r,n)$ are recalled in Sections 3.1 and 3.2, and the abstract Eqs. (8)-(10) are rewritten in a more concrete form in the special case when $\mathcal{M}={\rm Gr}_{\mathbb{C}}(r,n)$ in Sections 3.3 and 3.4. More details about Grassmann manifold can be found in the review articles in Refs. 49, 50 and the monographs in Refs. 51, 52.
@@ 3.1 Geometry of complex Grassmann manifolds
which in essence coincides with the exponential parametrization of molecular coefficients widely used in quantum chemistry 53.
@@ 3.2 Hamiltonian dynamics on Grassmann manifolds
[…]
@@
[…]
@@ 3.3 Linear energy functionals on projective spaces (FCI)
    a non-zero vector $\widetilde{\psi}\in{\mathbb{C}}^{n}$, it being understood that $\widetilde{\psi}$ and $\lambda\widetilde{\psi}$, with any $\lambda\in{\mathbb{C}}\setminus\{0\}$, represent the same element of ${\rm CP}^{n-1}$;
@@
$[H,\Gamma_{\star}]=0,\qquad\left\{\begin{array}[]{l}H\psi_{\star}=E\psi_{\star},\\ \|\psi_{\star}\|=1,\\ \psi_{\theta}(t)=e^{i\theta(t)}\psi_{\star},\end{array}\right.,\qquad\left\{\begin{array}[]{l}H\widetilde{\psi}_{\star}=E\widetilde{\psi}_{\star},\\ \widetilde{\psi}_{\lambda}(t)=e^{\lambda(t)}\psi_{\star}.\end{array}\right.$
@@
$-i[H,Q]=\left(\begin{array}[]{cc}0&((H|_{\Psi_{\star}^{\perp}}-E_{0})\Psi)^{*}\\ (H|_{\Psi_{\star}^{\perp}}-E_{0})\Psi&0\end{array}\right),$
@@
$i\frac{d\Psi}{dt}(t)=(H|_{\Psi_{\star}^{\perp}}-E_{0})\Psi(t)\quad\mbox{on}\quad\Psi_{\star}^{\perp},$
@@ 3.4 Mean-field models (TDHF and adiabatic TDDFT)
[…]
@@
$[J(\gamma)]_{pq}:=\sum_{r,s=1}^{N_{\rm b}}(pq|rs)\gamma_{sr},\quad[K(\gamma)]_{pq}:=\sum_{\mu,\nu=1}^{N_{\rm b}}(ps|rq)\gamma_{sr}.$ (34)
It follows that $H_{\gamma}:=\nabla E^{\rm GHF}(\gamma)=h+G(\gamma)$ is the Fock matrix, that (27) coincides with the usual TDHF equations, and (28) with the McWeeny formulation of the HF equations. In GHF theory, the no-unfilled shell property proved in Ref. 54 in the continuous setting also holds in the discretized setting. We therefore have that in the chosen basis of occupied/virtual real molecular spin-orbitals
@@
$\left(u_{1},\ldots,u_{N(N_{\rm b}-N)},J_{\gamma_{\star}}u_{1},\ldots,J_{\gamma_{\star}}u_{N(N_{\rm b}-N)}\right)$
of $T_{\gamma_{\star}}\mathcal{M}_{\mathbb{R}}$, build the matrix $\mathfrak{H}_{\star}$ defined in Eq. (11) with entries given by Eqs. (12)-(15), and compute the symplectic eigenvalues of $\mathfrak{H}_{\star}$. Choosing the canonical basis of $T_{\gamma_{\star}}\mathcal{M}_{\mathbb{R}}$, that is
$u_{[ia]}:=\frac{1}{\sqrt{2}}\left(\begin{array}[]{cc}0&e_{[ia]}\\ e_{[ia]}^{T}&0\end{array}\right),\quad J_{\gamma_{\star}}u_{[ia]}=\frac{1}{\sqrt{2}}\left(\begin{array}[]{cc}0&ie_{[ia]}\\ -ie_{[ia]}^{T}&0\end{array}\right),\quad\mbox{with}\quad[e_{[ia]}]_{jb}=\delta_{(ia),(jb)},$
for $1\leq i\leq N$ and $N+1\leq j\leq N_{\rm b}$, we deduce from Eq. (30), the fact that $D^{2}E^{\rm GHF}(\gamma)(Q)=G(Q)$, and the real character of the spin-orbitals that
$\mathfrak{H}_{\star}^{\rm GHF}=\left(\begin{array}[]{cc}A+B&0\\ 0&A-B\end{array}\right),$ (41)
@@
$A_{ai,bj}:=(\varepsilon_{a}-\varepsilon_{i})\delta_{ij}\delta_{ab}+(ai||jb),\quad B_{ai,bj}:=(ai||bj).$
@@
an expression in agreement with Ref. 55 Eqs. (II.78)-(II.79).
@@ 4 Theoretical analysis in the weakly-interacting regime
In this part, we aim to compare theoretically, in the weakly-interacting regime, the CP and LR formalisms for computing the electronic excitation energies at the FCI and HF levels of theory. We consider a family of Hamiltonians parametrized by an additional coupling parameter $\lambda$ that modulates the degree of interaction within the energy functional. In the non-interacting limit ($\lambda=0$), HF and FCI lead to the same ground and excited state energies, so that CP-HF provides the same sequence of excitation energies as FCI, while LR-HF returns the subsequence of excitation energies corresponding to single excitations. Under a nondegeneracy assumption, we are able to compute the FCI (Section 4.1), as well as CP-HF and LR-HF (Section 4.2) excitation energies up to second-order errors in $\lambda$ in the weakly-interacting regime ($|\lambda|\ll 1$). To our knowledge, this is the first study aiming at comparing the CP and LR methodologies from a theoretical point of view. These theoretical findings are substantiated by numerical results in Section 5. We shall remark that although the cases of $\lambda\neq 1$ are non-physical, our analysis could offer potential insights into the study of more complicated real molecules.
@@
$\widehat{H}_{\lambda}:=\widehat{h}+\lambda\widehat{V}=\sum_{\kappa,\lambda=1}^{N_{\rm b}}h_{\kappa\lambda}\sum_{\sigma\in\{\alpha,\beta\}}\widehat{a}_{\kappa\sigma}^{\dagger}\widehat{a}_{\lambda\sigma}+\frac{\lambda}{2}\sum_{\kappa,\lambda,\mu,\nu=1}^{N_{\rm b}}\sum_{\sigma,\sigma^{\prime}\in\{\alpha,\beta\}}(\kappa\sigma,\mu\sigma|\lambda\sigma^{\prime},\nu\sigma^{\prime})\widehat{a}_{\kappa\sigma}^{\dagger}\widehat{a}_{\lambda\sigma^{\prime}}^{\dagger}\widehat{a}_{\nu\sigma^{\prime}}\widehat{a}_{\mu\sigma},$
where $\lambda\geq 0$ is a coupling parameter. Let $(\phi_{n}^{0})_{1\leq n\leq N_{\rm b}}$ be an orthonormal basis of ${\mathbb{C}}^{N_{\rm b}}$ diagonalizing the one-electron Hermitian matrix $h$, such that
@@ 4.1 Analysis at the FCI level of theory
$\Phi_{0}^{(0)}:=\frac{1}{\sqrt{N!}}\;(\phi_{1}^{0}\otimes\alpha)\wedge(\phi_{1}^{0}\otimes\beta)\wedge\cdots\wedge(\phi_{N_{\rm p}}^{0}\otimes\alpha)\wedge(\phi_{N_{\rm p}}^{0}\otimes\beta),$
@@
$\Phi^{(0)}_{\alpha}$ $:=\frac{1}{\sqrt{N!}}\;(\phi_{1}^{0}\otimes\alpha)\wedge(\phi_{1}^{0}\otimes\beta)\wedge\cdots\wedge(\phi_{N_{\rm p}-1}^{0}\otimes\alpha)\wedge(\phi_{N_{\rm p}-1}^{0}\otimes\beta)\wedge(\phi_{N_{\rm p}+1}^{0}\otimes\alpha)\wedge(\phi_{N_{\rm p}}^{0}\otimes\beta),$
$\Phi^{(0)}_{\beta}$ $:=\frac{1}{\sqrt{N!}}\;(\phi_{1}^{0}\otimes\alpha)\wedge(\phi_{1}^{0}\otimes\beta)\wedge\cdots\wedge(\phi_{N_{\rm p}-1}^{0}\otimes\alpha)\wedge(\phi_{N_{\rm p}-1}^{0}\otimes\beta)\wedge(\phi_{N_{\rm p}}^{0}\otimes\alpha)\wedge(\phi_{N_{\rm p}+1}^{0}\otimes\beta).$
Note that $\Phi^{(0)}_{\alpha}$ and $\Phi^{(0)}_{\beta}$ are not eigenfunctions of $\bm{S}^{2}$, but that $\Psi_{\rm S}^{(0)}:=(\Phi^{(0)}_{\alpha}-\Phi^{(0)}_{\beta})/\sqrt{2}$ is a singlet state (the first-excited state in the $\bm{S}^{2}=0$ sector) and $\Psi_{\rm T}^{(0)}:=(\Phi^{(0)}_{\alpha}+\Phi^{(0)}_{\beta})/\sqrt{2}$ the $S_{z}=0$ component of a triplet state (the ground state in the $\bm{S}^{2}=2$ sector). The energies of the ground state $\Phi^{(0)}_{0}$ and the two-fold degenerate first excited state are respectively
@@
It follows from standard nondegenerate perturbation theory that for small $\lambda$, the ground state of $\widehat{H}_{\lambda}$ is nondegenerate and its energy $E_{0}(\lambda)$ is a real-analytic function of $\lambda$. Likewise the first excited state in the $\bm{S}^{2}=0$ sector as well as the ground state in the $\bm{S}^{2}=2$ sector are nondegenerate, and their energies $E_{\rm S}(\lambda)$ and $E_{\rm T}(\lambda)$ are real-analytic functions of $\lambda$ for $|\lambda|\ll 1$. Consequently, so are the FCI excitation energies $\omega_{\rm S}^{\rm FCI}(\lambda):=E_{\rm S}(\lambda)-E_{0}(\lambda)$ and $\omega_{\rm T}^{\rm FCI}(\lambda):=E_{\rm T}(\lambda)-E_{0}(\lambda)$:
$\omega_{\rm S}^{\rm FCI}(\lambda)=\sum_{k=0}^{+\infty}\lambda^{k}\omega^{{\rm FCI},(k)}_{\rm S},\quad\omega_{\rm T}^{\rm FCI}(\lambda)=\sum_{k=0}^{+\infty}\lambda^{k}\omega^{{\rm FCI}(k)}_{\rm T},$
@@
$\omega^{{\rm FCI},(0)}_{\rm S}=\omega^{{\rm FCI},(0)}_{\rm T}=\omega^{(0)}_{\rm HOMO-LUMO}:=\varepsilon_{N_{\rm p}+1}-\varepsilon_{N_{\rm p}}.$
@@
$\omega^{{\rm FCI},(1)}_{\rm S/T}=$ $\sum_{j=1}^{N_{\rm p}-1}((N_{\rm p}+1)\beta,(N_{\rm p}+1)\beta||j\beta,j\beta)-\sum_{j=1}^{N_{\rm p}-1}(N_{\rm p}\beta,N_{\rm p}\beta||j\beta,j\beta)$
$+\sum_{j=1}^{N_{\rm p}}((N_{\rm p}+1)\beta,(N_{\rm p}+1)\beta|j\alpha,j\alpha)-\sum_{j=1}^{N_{\rm p}}(N_{\rm p}\beta,N_{\rm p}\beta|j\alpha,j\alpha)$
±$\pm(N_{\rm p}\beta,(N_{\rm p}+1)\beta|(N_{\rm p}+1)\alpha,N_{\rm p}\alpha).$ (42)

### v2

@@ 3 Dynamics on Grassmann manifolds (FCI, TDHF, TDDFT)
where $r,n$ are positive integers such that $1\leq r\leq n-1$ (the trivial cases $\mathcal{M}={\rm Gr}_{\mathbb{C}}(0,n)=\{0\}$ and $\mathcal{M}={\rm Gr}_{\mathbb{C}}(n,n)=\{I_{n}\}$ are uninteresting). Such manifolds appear in quantum dynamics in at least two settings:
@@
    pure-state dynamics of generic isolated finite-dimensional quantum systems described by the linear Schrödinger equation, or equivalently the linear quantum Liouville equation (see Section 3.3). In computational quantum chemistry, this corresponds to the FCI level of theory;[35, 36]
@@
    mean-field models for many-body fermionic systems. This framework encompasses the cases of both TDHF [9, 10, 11, 12, 13, 14, 15] and adiabatic TDDFT,[20, 21, 22, 23, 24, 25] for which the dynamics is governed by a system of coupled nonlinear Schrödinger equations, or equivalently a nonlinear quantum Liouville equation (see Section 3.4).
Before elaborating on these two fundamental examples, the geometry and Kähler structure of the complex Grassmann manifold ${\rm Gr}_{\mathbb{C}}(r,n)$ are recalled in Sections 3.1 and 3.2, and the abstract Eqs. (9)-(12) are rewritten in a more concrete form in the special case when $\mathcal{M}={\rm Gr}_{\mathbb{C}}(r,n)$ in Sections 3.3 and 3.4. More details about Grassmann manifold can be found in the review articles [63, 64] and the monographs [65, 66].
@@ 3.1 Geometry of complex Grassmann manifolds
which in essence coincides with the exponential parametrization of molecular coefficients widely used in quantum chemistry. [67]
@@ 3.2 Hamiltonian dynamics on Grassmann manifolds
${[H_{\gamma_{\star}},\gamma_{\star}]=0}\quad\mbox{(critical points on Grassmann manifolds)}.$ (30)
@@
$\frac{dQ}{dt}(t)=-i[H_{\gamma_{\star}},Q]-i[D^{2}E(\gamma_{\star})Q,\gamma_{\star}]\quad\mbox{(linear response on Grassmann manifolds)}$} (33)
@@ 3.3 Linear energy functionals on projective spaces (FCI)
    a non-zero vector $\widetilde{\psi}\in{\mathbb{C}}^{n}$, it being understood that $\widetilde{\psi}$ and $\eta\widetilde{\psi}$, with any $\eta\in{\mathbb{C}}\setminus\{0\}$, represent the same element of ${\rm CP}^{n-1}$;
@@
$H,\Gamma_{\star}]=0,\qquad\left\{\begin{array}[]{l}H\psi_{\star}=E\psi_{\star},\\ \|\psi_{\star}\|=1,\\ \psi_{\theta}(t)=e^{i\theta(t)}\psi_{\star},\end{array}\right.,\qquad\left\{\begin{array}[]{l}H\widetilde{\psi}_{\star}=E\widetilde{\psi}_{\star},\\ \widetilde{\psi}_{\eta}(t)=e^{\eta(t)}\psi_{\star}.\end{array}\right$.
@@
$-i[H,Q]=\left(\begin{array}[]{cc}0&(-i(H|_{\Psi_{\star}^{\perp}}-E_{0})\Psi)^{*}\\ -i(H|_{\Psi_{\star}^{\perp}}-E_{0})\Psi&0\end{array}\right),$
@@
$\frac{d\Psi}{dt}(t)=-i(H|_{\Psi_{\star}^{\perp}}-E_{0})\Psi(t)\quad\mbox{on}\quad\Psi_{\star}^{\perp},$
@@ 3.4 Mean-field models (TDHF and adiabatic TDDFT)
$h_{pq}:=\int_{{\mathbb{R}}^{3}\times\{\alpha,\beta\}}\overline{\varphi_{p}(\bm{x})}\,\widehat{h}\varphi_{q}(\bm{x})\,{\rm d}\bm{x},\quad\text{where}~\widehat{h}~\text{is the one-body Hamiltonian,}$
@@
$J(\gamma)]_{pq}:=\sum_{r,s=1}^{N_{\rm b}}(pq|rs)\gamma_{sr},\quad[K(\gamma)]_{pq}:=\sum_{r,s=1}^{N_{\rm b}}(ps|rq)\gamma_{sr}.$ (36)
It follows that $H_{\gamma}:=\nabla E^{\rm GHF}(\gamma)=h+G(\gamma)$ is the Fock matrix, that (29) coincides with the usual TDHF equations, and (30) with the McWeeny formulation of the HF equations. In GHF theory, the no-unfilled shell property proved in Ref. [68] in the continuous setting also holds in the discretized setting. We therefore have that in the chosen basis of occupied/virtual real molecular spin-orbitals
@@
$\left(u_{1},\ldots,u_{N(N_{\rm b}-N)},-J_{\gamma_{\star}}u_{1},\ldots,-J_{\gamma_{\star}}u_{N(N_{\rm b}-N)}\right)$
of $T_{\gamma_{\star}}\mathcal{M}_{\mathbb{R}}$, build the matrix defined in Eq. (13) with entries given by Eqs. (14)-(17), and compute its symplectic eigenvalues. Choosing the canonical basis of $T_{\gamma_{\star}}\mathcal{M}_{\mathbb{R}}$, that is
$u_{[ia]}:=\frac{1}{\sqrt{2}}\left(\begin{array}[]{cc}0&e_{[ia]}\\ e_{[ia]}^{T}&0\end{array}\right),~~J_{\gamma_{\star}}u_{[ia]}=\frac{1}{\sqrt{2}}\left(\begin{array}[]{cc}0&ie_{[ia]}\\ -ie_{[ia]}^{T}&0\end{array}\right),~~\mbox{with}~~[e_{[ia]}]_{jb}=\delta_{(ia),(jb)},$
for $1\leq i,j\leq N$ and $N+1\leq a,b\leq N_{\rm b}$, we deduce from Eq. (32), the fact that $D^{2}E^{\rm GHF}(\gamma)(Q)=G(Q)$, and the real character of the spin-orbitals that
$\mathfrak{H}_{\star}^{\rm LR-GHF}=\left(\begin{array}[]{cc}A+B&0\\ 0&A-B\end{array}\right),$ (43)
@@
$A_{ia,jb}:=(\varepsilon_{a}-\varepsilon_{i})\delta_{ij}\delta_{ab}+(ai||jb),\quad B_{ia,jb}:=(ai||bj).$
@@
an expression in agreement with Ref. [69, Eqs. (II.78)-(II.79)].
Note that a more intrinsic way to write down the canonical basis of $T_{\gamma_{\star}}\mathcal{M}_{\mathbb{R}}$ is
$H_{\gamma_{\star}}=\sum_{n=1}^{N_{\rm b}}\varepsilon_{n}|\varphi_{n}\rangle\langle\varphi_{n}|,\quad\langle\varphi_{m}|\varphi_{n}\rangle=\delta_{mn},\quad\varepsilon_{1}\leq\cdots\leq\varepsilon_{N}<\varepsilon_{N+1}\leq\cdots\leq\varepsilon_{N_{\rm b}},$
$\gamma_{\star}=\sum_{i=1}^{N}|\varphi_{i}\rangle\langle\varphi_{i}|,\qquad T_{\gamma_{\star}}\mathcal{M}_{\mathbb{R}}={\rm Span}_{\mathbb{R}}\left(E_{ia},J_{\gamma_{\star}}E_{ia}\right),$
with
$E_{ia}:=\frac{|\varphi_{i}\rangle\langle\varphi_{a}|+|\varphi_{a}\rangle\langle\varphi_{i}|}{\sqrt{2}}\qquad\mbox{and}\qquad J_{\gamma_{\star}}E_{ia}=\frac{i|\varphi_{i}\rangle\langle\varphi_{a}|-i|\varphi_{a}\rangle\langle\varphi_{i}|}{\sqrt{2}}.$
@@ 4 Theoretical analysis in the weakly-interacting regime
In this part, we aim to compare theoretically, in the weakly-interacting regime, the CP and LR formalisms for computing the electronic excitation energies at the FCI and HF levels of theory. We consider a family of Hamiltonians parametrized by a coupling parameter $\eta$ that modulates the degree of interaction within the energy functional. In the non-interacting limit ($\eta=0$), HF and FCI lead to the same ground and excited state energies, so that CP-HF provides the same sequence of excitation energies as FCI, while LR-HF returns the subsequence of excitation energies corresponding to single excitations. Under a nondegeneracy assumption, we are able to compute the FCI (Section 4.1), as well as CP-HF and LR-HF (Section 4.2) excitation energies up to second-order errors in $\eta$ in the weakly-interacting regime ($|\eta|\ll 1$). To our knowledge, this is the first study aiming at comparing the CP and LR methodologies from a theoretical point of view. These theoretical findings are substantiated by numerical results in Section 5. We shall remark that although the cases of $\eta\neq 1$ are non-physical, our analysis could offer potential insights into the study of more complicated real molecules.
@@
$\widehat{H}_{\eta}:=\widehat{h}+\eta\widehat{V}=\sum_{\kappa,\lambda=1}^{N_{\rm b}}h_{\kappa\lambda}\sum_{\sigma\in\{\alpha,\beta\}}\widehat{a}_{\kappa\sigma}^{\dagger}\widehat{a}_{\lambda\sigma}+\frac{\eta}{2}\sum_{\kappa,\lambda,\mu,\nu=1}^{N_{\rm b}}\sum_{\sigma,\sigma^{\prime}\in\{\alpha,\beta\}}(\kappa\mu|\lambda\nu)_{\sigma\sigma^{\prime}}\widehat{a}_{\kappa\sigma}^{\dagger}\widehat{a}_{\lambda\sigma^{\prime}}^{\dagger}\widehat{a}_{\nu\sigma^{\prime}}\widehat{a}_{\mu\sigma},$
where $\eta\geq 0$ is a coupling parameter and the spin dependence of the integrals is indicated with the notation
$(\kappa\mu||\lambda\nu)_{\sigma\sigma}:$ $=(\kappa_{\sigma}\mu_{\sigma}|\lambda_{\sigma}\nu_{\sigma})-(\kappa_{\sigma}\nu_{\sigma}|\lambda_{\sigma}\mu_{\sigma}),~\sigma\in\{\alpha,\beta\},$
$(\kappa\mu|\lambda\nu)_{\sigma\sigma^{\prime}}:$ $=(\kappa_{\sigma}\lambda_{\sigma}|\mu_{\sigma^{\prime}}\nu_{\sigma^{\prime}}),~\sigma\neq\sigma^{\prime}\in\{\alpha,\beta\}.$
Let $(\phi_{n}^{0})_{1\leq n\leq N_{\rm b}}$ be an orthonormal basis of ${\mathbb{C}}^{N_{\rm b}}$ diagonalizing the one-electron Hermitian matrix $h$, such that
@@ 4.1 Analysis at the FCI level of theory
$\Phi_{0}^{(0)}:=\frac{1}{\sqrt{N!}}\;\phi_{1}^{\alpha,0}\wedge\phi_{1}^{\beta,0}\wedge\cdots\wedge\phi_{N_{\rm p}}^{\alpha,0}\wedge\phi_{N_{\rm p}}^{\beta,0},$ (44)
where we adopt the notations
$\phi_{n}^{\sigma,0}:=\phi_{n}^{0}\otimes\sigma,\quad n=1,\ldots,N_{\rm b},\quad\sigma\in\{\alpha,\beta\},$
@@
$\Phi^{(0)}_{\alpha}$ $:=\frac{1}{\sqrt{N!}}\;\phi_{1}^{\alpha,0}\wedge\phi_{1}^{\beta,0}\wedge\cdots\wedge\phi_{N_{\rm p}-1}^{\alpha,0}\wedge\phi_{N_{\rm p}-1}^{\beta,0}\wedge\phi_{N_{\rm p}+1}^{\alpha,0}\wedge\phi_{N_{\rm p}}^{\beta,0},$
$\Phi^{(0)}_{\beta}$ $:=\frac{1}{\sqrt{N!}}\;\phi_{1}^{\alpha,0}\wedge\phi_{1}^{\beta,0}\wedge\cdots\wedge\phi_{N_{\rm p}-1}^{\alpha,0}\wedge\phi_{N_{\rm p}-1}^{\beta,0}\wedge\phi_{N_{\rm p}}^{\alpha,0}\wedge\phi_{N_{\rm p}+1}^{\beta,0}.$
Note that $\Phi^{(0)}_{\alpha}$ and $\Phi^{(0)}_{\beta}$ are not eigenfunctions of $\hat{S}^{2}$, but that $\Psi_{\rm S}^{(0)}:=(\Phi^{(0)}_{\alpha}-\Phi^{(0)}_{\beta})/\sqrt{2}$ is a singlet state (the first-excited state in the $\hat{S}^{2}=0$ sector) and $\Psi_{\rm T}^{(0)}:=(\Phi^{(0)}_{\alpha}+\Phi^{(0)}_{\beta})/\sqrt{2}$ the $S_{z}=0$ component of a triplet state (the ground state in the $\hat{S}^{2}=3/4$ sector). The energies of the ground state $\Phi^{(0)}_{0}$ and the two-fold degenerate first excited state are respectively
@@
It follows from standard nondegenerate perturbation theory that for small $\eta$, the ground state of $\widehat{H}_{\eta}$ is nondegenerate and its energy $E_{0}(\eta)$ is a real-analytic function of $\eta$. Likewise the first excited state in the $\hat{S}^{2}=0$ sector as well as the ground state in the $\hat{S}^{2}=3/4$ sector are nondegenerate, and their energies $E_{\rm S}(\eta)$ and $E_{\rm T}(\eta)$ are real-analytic functions of $\eta$ for $|\eta|\ll 1$. Consequently, so are the FCI excitation energies $\omega_{\rm S}^{\rm FCI}(\eta):=E_{\rm S}(\eta)-E_{0}(\eta)$ and $\omega_{\rm T}^{\rm FCI}(\eta):=E_{\rm T}(\eta)-E_{0}(\eta)$:
$\omega_{\rm S}^{\rm FCI}(\eta)=\sum_{k=0}^{+\infty}\eta^{k}\omega^{{\rm FCI},(k)}_{\rm S},\quad\omega_{\rm T}^{\rm FCI}(\eta)=\sum_{k=0}^{+\infty}\eta^{k}\omega^{{\rm FCI},(k)}_{\rm T}$,
@@
$\omega^{{\rm FCI},(0)}_{\rm S}=\omega^{{\rm FCI},(0)}_{\rm T}=\varepsilon_{N_{\rm p}+1}-\varepsilon_{N_{\rm p}}.$
@@
$\omega^{{\rm FCI},(1)}_{\rm S/T}$= $\sum_{j=1}^{N_{\rm p}-1}(N_{\rm p}+1,N_{\rm p}+1||j,j)_{\beta\beta}-\sum_{j=1}^{N_{\rm p}-1}(N_{\rm p},N_{\rm p}||j,j)_{\beta\beta}$
$+\sum_{j=1}^{N_{\rm p}}(N_{\rm p}+1,N_{\rm p}+1|j,j)_{\beta\alpha}-\sum_{j=1}^{N_{\rm p}}(N_{\rm p},N_{\rm p}|j,j)_{\beta\alpha}$
$\pm(N_{\rm p},N_{\rm p}+1|N_{\rm p}+1,N_{\rm p})_{\beta\alpha}.$ (45)
The first-order variation of the lowest excitation energy in the weakly-interacting regime ($0<\eta\ll 1$) is thus given by the triplet state.